OpenStudy (anonymous):

\(p\) is a an odd prime number.prove that \( 1^{p-1} + 2^{p-1} + 3^{p-1} + ... + (p-1)^{p-1} \equiv p + (p-1)! ~~ (mod ~ p^2) \)

OpenStudy (michele_laino):

I'm very sorry, I don't know your answer since I'm not an expert in discrete mathematics

OpenStudy (anonymous):

no prob ;) i just asked it for fun ;)

ganeshie8 (ganeshie8):

Wilson theorem!

OpenStudy (kainui):

My intuition is telling me that p and (p-1)! come from the binomial theorem, and the reason it's prime is so mod (p^2) has no other weird divisors to cut it up possibly. Other than that, I know how to derive individual results for sum of squares, cubes, etc... but not really sure how to take it from here yet, but maybe this helps someone out or they can correct me.

OpenStudy (kainui):

Here's some math that might get us to something to play with: \[\Large \sum_{n=1}^{p-1}(n+1)^{p-1}-n^{p-1}=p^{p-1}-1\] start with this telescoping series. Apply the binomial theorem on the first term in the sum and then immediately pull out the first term to cancel out with the other term leaving us with: \[\Large \sum_{n=1}^{p-1}\sum_{k=1}^{p-1}\left(\begin{matrix}p-1 \\ k\end{matrix}\right)n^{p-1-k} = p^{p-1}-1 \] Now I'm thinking where to go from here.

OpenStudy (kainui):

\[\Large \sum_{n=1}^{p-1}n^{p-1} = p+(p-1)! \mod (p^2)\] Is closely related to what I have up there, but it needs some tweaking to make it fit... bleh I don't know lol.

OpenStudy (kainui):

Alright I'm looking up wilson theorem I give up lol

OpenStudy (kainui):

Alternative statement, tried to do the gauss method lol. \[\Large \frac{1}{2} \sum_{n=1}^{\frac{p-1}{2}} \left[ (p-n)^{p-1}+n^{p-1} \right] \equiv p+(p-1)! \mod(p^2) \]

OpenStudy (anonymous):

@ganeshie8 it's mod p^2..wilson thm would not seem useful. guys a small help:u can cleverly use Fermat's little thm ;)

OpenStudy (anonymous):

consider that \(i^{p-1} \equiv 1 ~~ (mod ~ p)\) so \(\Large \frac{ i^{p-1}-1 }{ p }\) is an integer.

OpenStudy (anonymous):

are u using this ? \(1^{p-1}=1 \mod p\) \(1^{p-1}=1 \mod p^2\)

OpenStudy (anonymous):

hmm im trying this \(a^{\phi (p^2)}=1 \mod p^2\\ a^{p^2-p}=1 \mod p^2\\ a^{p-1}a^p=1 \mod p^2\\ \)

OpenStudy (anonymous):

not a bad idea,but as i said,if we consider that an integer,then we can prove it.

OpenStudy (anonymous):

\(\large \phi(p^2) = p(p-1)\) would help?

OpenStudy (anonymous):

btw i wish @ganeshie8 was here working with him makes it more fun , it remindes me with Wilson as well

OpenStudy (anonymous):

yes we can use wilson thm but at the finishing of the proof! u want to see a part of the proof?

ganeshie8 (ganeshie8):

By wilson \[(p-1)! \equiv -1\pmod p\] multiplying \(p\) through out gives \[p(p-1)! \equiv -p\pmod {p^2}\] adding \(p^2\) both sides \[p[p+(p-1)!] \equiv p(p-1)\pmod {p^2}\] fermat gives \[1^{p-1}+2^{p-1}+\cdots + (p-1)^{p-1}\equiv 1+1+\cdots + (p-1~\text{times})\\\equiv p-1 \pmod {p}\] \(\cdots\)

OpenStudy (anonymous):

i dont get it yet >.<

ganeshie8 (ganeshie8):

yeah thats not a proof, actually i got stuck after that :P

OpenStudy (anonymous):

hahaha

ganeshie8 (ganeshie8):

\[p[p+(p-1)!] \equiv p(p-1)\pmod {p^2}\] this should be easy to see right ?

OpenStudy (anonymous):

yes but how does it help

ganeshie8 (ganeshie8):

i got excited by noticing the p+(p-1)! term on left hand side

ganeshie8 (ganeshie8):

did u get why fermat gives \[1^{p-1}+2^{p-1}+\cdots + (p-1)^{p-1}\equiv 1+1+\cdots + (p-1~\text{times})\\\equiv p-1 \pmod {p}\] ?

OpenStudy (anonymous):

hmmm it should be p(p-1)/2 , right ?

ganeshie8 (ganeshie8):

\[a^{p-1}\equiv 1\pmod p \] since there are \(p-1\) terms in the sum, the sum evaluates to \(p-1\) in mod \(p\)

ganeshie8 (ganeshie8):

but its not easy to work \(p^2\) here

OpenStudy (anonymous):

ok lol got it xD i was seeing something else with that sum wew!

OpenStudy (anonymous):

keep the idea i gave you'll be able to prove it.

ganeshie8 (ganeshie8):

dont reveal the proof or hints

ganeshie8 (ganeshie8):

I end up with \[p[1^{p-1} + 2^{p-1} + 3^{p-1} + ... + (p-1)^{p-1}] \equiv p[p + (p-1)! ]\pmod {p^2}\] but we cannot cancel p just like that :/ ok im ready for hint if others are also ready..

ganeshie8 (ganeshie8):

@Marki

OpenStudy (anonymous):

hehe i died already xD

OpenStudy (anonymous):

i won't reveal ;) i'm eager to see ur ideas and gather them too :)

OpenStudy (anonymous):

well,consider that if we suppose \( r_{i} = \Large \frac{ i^{p-1}-1 }{ p } \) \( \prod_{i=1}^{p-1}i^{p-1} = \prod_{i=1}^{p-1}pr_{i}+1 = (pr_{1}+1)(pr_{2}+1)...(pr_{p-1}+1)\equiv 1+ p(r_1 + r_2 +... r_{p-1}) \)

OpenStudy (anonymous):

as \( p^2 | (pr_i)(pr_j) \) of the product of more than 2 of them

OpenStudy (anonymous):

u know and \( \prod_{i=1}^{p-1}i^{p-1} = (\prod_{i=1}^{p-1}i)^{p-1} \)

OpenStudy (anonymous):

now use wilson thm ;)

ganeshie8 (ganeshie8):

Amazing! okay that equals \((p-1)!^{p-1}\) and by wilson thm does that reduce to \(1\pmod p\) ?

ganeshie8 (ganeshie8):

and you have \[\prod_{i=1}^{p-1}i^{p-1} \equiv 1+ p(r_1 + r_2 +... r_{p-1})\pmod{p^2}\]

ganeshie8 (ganeshie8):

it how are you going to combine them xD

OpenStudy (anonymous):

nice ;) keep going.

ganeshie8 (ganeshie8):

i guess next step would be \[\sum \limits_{i=1}^{p-1}i^{p-1} \equiv \sum \limits_{i=1}^{p-1}(1+pr_i) \equiv p-1 + p(r_1 + r_2 +... r_{p-1})\pmod{p^2}\]

OpenStudy (anonymous):

correct ;)

ganeshie8 (ganeshie8):

then so far we have :\[\sum \limits_{i=1}^{p-1}i^{p-1} \equiv p-2+ (p-1)!^{p-1}\pmod{p^2}\]

ganeshie8 (ganeshie8):

that right hand side has to reduce to p + (p-1)! somehow interesting, i don't see how wilson helps here yet :O

OpenStudy (anonymous):

well,we can do the same thing as we did!how about write \( \Large \frac{(p-1)! +2}{p} = r\) and find \( (p-1)!^{p-1} \) from this and play with ?

OpenStudy (anonymous):

i give up :|

ganeshie8 (ganeshie8):

Oh yes i know you meant \[\frac{(p-1)! +1}{p} = r\] right ?

ganeshie8 (ganeshie8):

@Marki this is going to be fun xD we're almost done here

OpenStudy (anonymous):

sry yes ;)

ganeshie8 (ganeshie8):

\[(p-1)!^{p-1} \equiv (-1+pr)^{p-1} \equiv (pr-1)^{p-1} \equiv pr + 1 \pmod{p^2}\]

ganeshie8 (ganeshie8):

does that look okay ? i think il need to replace r

OpenStudy (anonymous):

awesome ;)

OpenStudy (anonymous):

but wait,it would be \( 1 - rp(p-1) \) not \( pr+1 \)

OpenStudy (anonymous):

:O

ganeshie8 (ganeshie8):

(Awesome+1)!^(Awesome) this is the coolest NT proof i have seen in like months

ganeshie8 (ganeshie8):

\(1-rp(p-1)\equiv pr+1 \pmod {p^2}\) @PFEH.1999

OpenStudy (anonymous):

oh correct ;)

OpenStudy (anonymous):

so it would be simply done ;)

ganeshie8 (ganeshie8):

pluggin back r we get \[\begin{align}(p-1)!^{p-1} \equiv (-1+pr)^{p-1} \equiv (pr-1)^{p-1} &\equiv pr + 1 \pmod{p^2}\\~\\ &\equiv (p-1)!+2 \pmod{p^2} \end{align}\]

ganeshie8 (ganeshie8):

finally \[\sum \limits_{i=1}^{p-1}i^{p-1} \equiv p-2+ (p-1)!^{p-1}\pmod{p^2}\] becomes \[\sum \limits_{i=1}^{p-1}i^{p-1} \equiv p + (p-1)!\pmod{p^2}\] XD

OpenStudy (anonymous):

nice ;) this problem is really awesome,don't u think so?

OpenStudy (anonymous):

yes i do :D

OpenStudy (anonymous):

well,i'll come with more exciting problems ;)

OpenStudy (anonymous):

so btw does this reduce the problem to -(p+1) mod p^2 ?

OpenStudy (anonymous):

what?!

ganeshie8 (ganeshie8):

wil look forward to attempt them @PFEH.1999 xD @Marki are u trying to reduce p + (p-1)! in mod p^2 ?

OpenStudy (anonymous):

like show that \(1^{p-1} + 2^{p-1} + 3^{p-1} + ... + (p-1)^{p-1} \equiv p + (p-1)!\equiv -(p+1) ~~ (mod ~ p^2)\)

ganeshie8 (ganeshie8):

i think below is trivially true by fermat's little thm : \[1^{p-1} + 2^{p-1} + 3^{p-1} + ... + (p-1)^{p-1} \equiv p-1 \pmod p\] but \(\mod p^2\) wont reduce further

OpenStudy (anonymous):

i had a feeling it would work

OpenStudy (anonymous):

well,not true XD \( p =5 \) so \( 5+ 24 \equiv -6 \) but 35 is not divisble by 25.

OpenStudy (anonymous):

haha i see :P

OpenStudy (anonymous):

@ganeshie8 , awesome site ;) i tried to give it a 20 digits number to tell me whether it's prime or not ;D

OpenStudy (anonymous):

lol @PFEH.1999 u havnt troll it enough :P

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

a guy told me if u want to improve ur mind before sleeping try to check all primes smaller than 2000 !

OpenStudy (anonymous):

i havn't tried so far ;) i'm afraid i might burn my mind XD

OpenStudy (anonymous):

interesting , i do similar stuff but mostly it distract me and keep me away from sleep

OpenStudy (anonymous):

try to find the happy numbers lol xD or weird numbers

OpenStudy (anonymous):

it would be more quick than primes

OpenStudy (anonymous):

i would like other question xD

OpenStudy (anonymous):

it's easier ;)

OpenStudy (anonymous):

ahaha xD its painful who said that counting sheep makes u sleep like if i tried to work on numbers before sleep that means no sleep at all